OpenStudy (anonymous):

Find the solution to the equation 64^(2 - x) = 4^4x. Using complete sentences, explain the procedure used to solve this equation

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

im sorry I forgot to add a couple things in my equiation 64^(2 - x) = 4^4x

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

\(64=4^3\) and so \(64^{2-x}=(4^3)^{2-x}=4^{6-3x}\)

OpenStudy (anonymous):

now your equation is \[4^{6-3x}=4^{4x}\] making \(6-3x=4x\) solve for \(x\)

OpenStudy (anonymous):

@satellite73 so I would end up with 7x=6 then x=6/7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if that is my answer how would I put that into a sentence? by the way thank you for your help! I appreciate it very much @satellite73